OpenStudy (anonymous):

12 - 6 + (14 - 8)to the 2nd power i need a numerical answer

OpenStudy (vlery):

144

OpenStudy (anonymous):

12-6+(14-8)^2 12-6+6^2 12-6+36 6+36 42

OpenStudy (vlery):

12-6=6 14-8=6 6+6=12 12^2=144

OpenStudy (anonymous):

thank you so much for your help but which one is it?

OpenStudy (vlery):

I think insta was correct sorry for the wrong answer

OpenStudy (anonymous):

thank you guys!